ULTRAVIOLET PHOTODISSOCIATION DYNAMICS OF THE 1-PROPENYL RADICAL

\underline{MICHAEL LUCAS}, Department of Chemistry, University of California at Riverside, Riverside, CA 92521; YU SONG, Department of Chemistry, University of California at Davis, Davis, CA 95616; JINGSONG ZHANG ^{a}, Department of Chemistry, University of California at Riverside, Riverside, CA 92521; CHRISTOPHER BRAZIER, Department of Chemistry and Biochemistry, California State University, Long Beach, Long Beach, CA 90840.

Ultraviolet (UV) photodissociation dynamics of jet-cooled 1-propenyl radical (C_{3}H_{5}) were studied in the photolysis wavelength region of 224 to 248 nm using high-n Rydberg atom time-of-flight (HRTOF) and resonance enhanced multiphoton ionization (REMPI) techniques. The 1-propenyl radical were produced from 193-nm photolysis of 1-chloropropene and 1-bromopropene precursors. The H-atom photofragment yield (PFY) spectra have a broad peak centered around 230 nm. The H + C_{3}H_{4} product translational energy distributions, P(E_{T})’s, peak near ∼ 10 kcal/mol, and the fraction of average translational energy in the total excess energy, ⟨f_{T} ⟩, is nearly a constant of ∼ 0.14 from 224 to 248 nm. The P(E_{T})’s along with recent theoretical calculations indicate that H + propyne dissociation is the predominant H-atom product channel. The H-atom product angular distribution is isotropic, with the anisotropy parameter β ∼ 0. The dissociation mechanism is consistent with unimolecular dissociation of hot 1-propenyl radical after internal conversion from the electronically excited state.

aCorresponding author.